3 ladies and 3 gents can be seated at a round table sothat any two and only two of ladies sit together number of ways is?
Answer by Edwin McCravy(20054)

When the problem says "round table", that means it's as if the table
and people were on a turntable floor.  So there are only these two
seating schemes where exactly 2 women sit together.


   W   W                W   W
  
 M       M            M       M
   
   M   W                W   M

For each of these two arrangments there are 3! ways to place the women
and 3! ways to place the men.  Thats 3!*3! = 6*6 = 36 ways for each.
Since there are 2 seating schemes above, we double that number.  

Answer = 2*36 = 72

Edwin